OpenStudy (anonymous):

HELP!!!! I put a pic up Evaluate the series E with 4 on top n-1 on bottom and (n+4) to the right. a:26 b:10 c:16 d:-6

OpenStudy (anonymous):

\[\sum_{n-1}^{4}(n+1)\]

OpenStudy (anonymous):

should be n+4 not n+1

OpenStudy (ranga):

Is it n-1 at the bottom of the sigma symbol or n = 1

OpenStudy (anonymous):

it's a minus sign

OpenStudy (ranga):

Can you put the picture of the problem up?

OpenStudy (anonymous):

I will try. I have to scan it in to my computer

OpenStudy (ranga):

Okay, I will wait for the problem to be posted.

OpenStudy (anonymous):

ok - if n=1 then I would just do 1+4=5, 2+4=6. 3+4-7,4+4=8 then add them together to get 26

OpenStudy (anonymous):

Is that correct?

OpenStudy (ranga):

I don't see this problem in the posted pic.

OpenStudy (ranga):

I think in the problems you have posted it is all "n = " and not "n -"

OpenStudy (ranga):

Yes 26 is the correct answer for this problem.

OpenStudy (ranga):

All the rest of the problems you do the same way. Put n = 1, 2, 3, etc. till the upper limit, evaluate the expression and add them all up.